• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 21:14:20 on February13,2009

    Entry number 260982


    keyword=Gain Shift in Compton Photon Detector

    Comparing measurements this week to measurements taken in December, we see a 20% gain shift in the Compton photon detector. The blue line shows the Compton spectrum measured in December at 12 uA; the red line shows the Compton spectrum measured today at 15 uA. Both measurements were taken at 5 pass with the GSO's PMT set at -1800 V.

    We suspect that this is related to a rate-dependent gain shift that we observed in tests at HIGS in the fall of 2008 (same crystal, same PMT model, similar tube base); the gain seems to stay relatively constant over several days. If this is the case, then it should not have a large effect on asymmetry measurements.

    FIGURE 1